my tens digit is 3 more than my ones digit. my hundreds is 1 less than my tens digit.what number am I if the sum of my digits is 20?
Answer by Edwin McCravy(20056)


>>...my tens digit is 3 more than my ones digit...<<
equation 1.   t = u+3

>>...my hundreds is 1 less than my tens digit...<<
equation 2.   h = t-1

>>...the sum of my digits is 20...<<
equation 3.   h+t+u = 20

Use equation 1 to substitute u+3 for t in equation 2:

equation 2.  h = t-1
             h = (u+3)-1
             h = u+3-1
equation 4.  h = u+2

Use equation 1 and equation 4 to substitute u+3 for t 
and u+2 for h in equation 3:

equation 3.           h+t+u = 20
              (u+2)+(u+3)+u = 20
                  u+2+u+3+u = 20
                       3u+5 = 20
                         3u = 15
equation 5.               u = 5

Use equation 5 to substitute 5 for u in equation 4

equation 4.  h = u+2
             h = 5+2
             h = 7

Use equation 5 to substitute 5 for u in equation 1

equation 1.   t = u+3
              t = 5+3
              t = 8

So the digits are h = 7, t = 8, and u = 5 and
therefore the number is 785.

Edwin